Project L costs $55,000, its expected cash inflows are $10,000 per year for 12 years, and its WACC is 11%. What is the project's payback? Round your answer to two decimal places. ____years
Answer #1

Answer:

Normal Payback period:

Given

Initial cost P=$55000

Annual Cash flow A=$10000

So since in 5 years we can recover $50000(5*10000=50000) so

pay back period = 5+(55000-50000)/1000-=5.5 years

Discounted payback period

Discounted cash flow = CF/(1+WACC)^t

CF= cash flow

WACC=11%

t=year in which cash flow occur

Year cash flow discounted cash flow Cumulative discounted cash flow
1 10000 9009.01 9009.01
2 10000 8116.22 17125.23
3 10000 7311.91 24437.15
4 10000 6587.31 31024.46
5 10000 5934.51 36958.97
6 10000 5346.41 42305.38
7 10000 4816.58 47121.96
8 10000 4339.26 51461.23
9 10000 3909.25 55370.48
10 10000 3521.84 58892.32
11 10000 3172.83 62065.15
12 10000 2858.41 64923.56

From above table we find that payback period will be between 8 and 9 so
Discounted pay back period = 8+(55000-51461.23)/3909.25=8.91 years